OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

OpenStudy (anonymous):

the equation in slope intercept form is y=(3/2)x + 2 The y intercept of this equation is 2, and it has the points (0,2) and (2,5) and (4, 8) In order for a line to be parallel, it has to have the same slope so automatically, the slope of the answer is also (3/2). When it is at x=4, y =2 looking at the other line, it shows when x=4, y=8 The difference is 6, vertically, so you just have to shift the line down 6 units so if its positive 2 as the y intercept at first, move it down six, making it -4 thus the equation is y=(3/2)x - 4

OpenStudy (anonymous):

thats wrong but thank you for helping (: